A collision is shown here (left side: before the collision and right side: after the collision). Use M1 = 2 kg, M2 = 4 kg, V1 = 3 m/s, V2 = ?1 m/s and V3 = 2 m/s. (a) Determine if this collision satisfies linear momentum conservation. (b) Determine if this collision is elastic.A collision is shown here (left side: before the c

Please provide all your work. Thank you.
Answer #1

here,

m1 = 2 kg

m2 = 4 kg

v1 = 3 m/s

v2 = -1 m/s

v3 = 2 m/s

(a)

initial momentum , Pi = m1*v1

Pi = 6 kg . m/s

final momentum , Pf = m1*v2 + M2*v3

Pf = 2*(-1) + 4 * 2


Pf = 6 kg .m/s

initial momentum is equal to final momentum

so it satisfy linear momentum conservation

(b)

initial kinetic energy , KEi = 0.5 * m1*v1^2

KEi = 9 J

final kinetic energy , KEf = 0.5 * m1*v2^2 + 0.5 * M2*v3^2

KEf = 0.5 * 2*(-1)^2 + 0.5 * 4 * 2^2


KEf = 9 J


initial kinetic energy is equal to final kinetic energy

so the collison is elastic